Find the indicated area under the standard normal curve. To the left of zequalsnegative 1.56 and to the right of zequals1.56 Click here to view page 1 of the standard normal table. LOADING... Click here to view page 2 of the standard normal table. LOADING... The total area to the left of zequalsnegative 1.56 and to the right of zequals1.56 under the standard normal curve is nothing. ​(Round to four decimal places as​ needed.)
Answer #1

Ans:

For standard normal curve:

P(z<-1.56 and z>1.56)=P(z<-1.56)+P(z>1.56)

=2*P(z<-1.56)

=2*0.0594

=0.1188